# An Efficient Algorithm for Generating Minimal Unique-Cause MC/DC Test cases for Singular Boolean Expressions

### v2

@@ Abstract.
ccs: Software and its engineering Software testing and debugging
@@ 3.1. Phase 1: Preprocessing
5:         $R\leftarrow\text{apply\_DeMorgan\_recursively}(S)$triangleright Phase 1.1: Apply De Morgan’s
6:         $R\leftarrow\text{sort\_conditions}(R)$triangleright Proceed to Phase 1.2
@@
8:         $R\leftarrow\text{sort\_conditions}(S)$triangleright Phase 1.2: Sort directly
@@ 3.3.1. Terminology and Notation
$RT$ Relation Table A truth value sequence derived from the SBE’s operator structure.
@@ 3.3.3. Relation Table Generation
Once the structural decomposition is complete, the algorithm analyzes the hierarchy of operators in the decomposed structure to generate a Relation Table (RT). The purpose of this table is to convert the complex, nested operator relationships into a simple, linear sequence of truth values that can be used in Phase 4. For instance, for a structure $B_{1}^{\wedge}\wedge B_{2}^{\vee}$, the algorithm identifies the operator combination (,$\wedge$,$\wedge$,$\vee$ , ∧ , ∨) and, according to predefined rules, converts it into a final truth value sequence, ‘RT(T, T, F)’. This table acts as a clear blueprint that dictates which base patterns to apply and in what order during the test case composition.
@@ 3.4. Phase 4: Test Case Generation
1:Input: $S_{norm}$ (A preprocessed SBE), $RT$ (The Relation Table from Phase 3).
@@
3:$SubExprs\leftarrow\textsc{DecomposeSBE}(S_{norm})$
4:$N\leftarrow\textsc{CountOperands}(S_{norm})$
5:$T_{final}\leftarrow\textsc{InitializeTable}(N+1,N)$
@@
7$\triangleright$ Iterate through SBE blocks, referencing pre-generated RT to expand the table
8:while $SubExprs$ is not empty do
@@
11$\triangleright$ Table combination and expansion using RT
12:     if $NextBlock$ is a B-Form then
@@
16:     else if $NextBlock$ is an S-Form then
[…]
18:         $\textsc{FillSColumn}(T_{expanded},RT,Op)$triangleright Use RT per Sec. 3.4.2
19:     end if
20$\triangleright$ Handle exceptional cases during composition using RT

### v3

@@ Abstract.
$c$s: Software and its engineering Software testing and debugging
@@ 3.1. Phase 1: Preprocessing
5:   $R\leftarrow\text{apply\_DeMorgan\_recursively}(S)$ $\triangleright$ Phase 1.1: Apply De Morgan’s
6:   $R\leftarrow\text{sort\_conditions}(R)$ $\triangleright$ Proceed to Phase 1.2
@@
8:   $R\leftarrow\text{sort\_conditions}(S)$ $\triangleright$ Phase 1.2: Sort directly
@@ 3.3.1. Terminology and Notation
RTRT Relation Table A truth value sequence derived from the SBE’s operator structure.
@@ 3.3.3. Relation Table Generation
Once the structural decomposition is complete, the algorithm analyzes the hierarchy of operators in the decomposed structure to generate a Relation Table (RT). The purpose of this table is to convert the complex, nested operator relationships into a simple, linear sequence of truth values that can be used in Phase 4. For instance, for a structure $B_{1}^{\wedge}\wedge B_{2}^{\vee}$, the algorithm identifies the operator combination ($\wedge,\wedge,\vee$) and, according to predefined rules, converts it into a final truth value sequence, ‘RT(T, T, F)’. This table acts as a clear blueprint that dictates which base patterns to apply and in what order during the test case composition.
@@ 3.4. Phase 4: Test Case Generation
1:Input: $S_{norm}$ (A preprocessed SBE), RTRT (The Relation Table from Phase 3).
@@
[…]
8:while SubExprsSubExprs is not empty do
@@
[…]
12:  if NextBlockNextBlock is a B-Form then
@@
15:   $\textsc{FillEmptyRegions}(T_{expanded},RT)$ $\triangleright$ Use RT per Sec. 3.4.1
16:  else if NextBlockNextBlock is an S-Form then
17:   $T_{expanded}\leftarrow\textsc{LayoutTables}(T_{current},\text{S\_Column})$
18:   $\textsc{FillSColumn}(T_{expanded},RT,Op)$ $\triangleright$ Use RT per Sec. 3.4.2
19:  end if